# TARS [Uncaptioned image]: MinMax Token-Adaptive Preference Strategy for MLLM Hallucination Reduction

### v3

@@ Appendix D Additional Experimental Results
We present extended results on the AMBER benchmark in Table 6, evaluating hallucination performance from both generative and discriminative perspectives.
@@
Beyond generative evaluation, we further introduce fine-grained discriminative metrics that assess hallucination across four categories: Existence, Relation, Attribute, and Action. This allows a more detailed understanding of where hallucinations occur and how well each method suppresses them. As shown in the right half of the table, TARS consistently outperforms both DPO and the LLaVA baseline in all dimensions. Notably, it excels in Relation and Attribute grounding, where conventional methods often struggle due to subtle cross-modal mismatches.
Together, these results underscore the strength of our token-adaptive perturbation strategy, which not only reduces hallucinations at a global level but also enhances semantic fidelity in specific visual grounding aspects, without relying on hand-crafted heuristics or additional supervision.
To further dissect model performance on hallucination-prone scenarios, we report average scores across different question categories in the MMHal benchmark (Figure 4). TARS consistently achieves higher scores across all categories, particularly excelling in fine-grained tasks involving spatial reasoning and attribute identification. These results suggest that our method improves not only overall hallucination rates but also robustness to diverse multimodal challenges, highlighting its effectiveness in aligning responses with nuanced visual semantics.
Table 6: Comparison of generative and fine-grained discriminative hallucination metrics on the AMBER benchmark. TARS achieves consistent gains over DPO and the LLaVA baseline across both holistic and category-specific evaluations, demonstrating enhanced visual-textual alignment and robust hallucination suppression.
@@
LLaVA-v1.5-7B (Liu et al., 2023b) 7.9 54.7 37.1 3.2 82.9 58.6 65.6 70.1
+ DPO (Li et al., 2024) 4.9 56.6 26.4 2.5 87.1 59.7 74.6 79.4
+ TARS (Ours) 2.4 59.6 13.2 0.4 95.3 62.8 78.6 86.5
LLaVA-v1.5-13B (Liu et al., 2023b) 6.7 52.1 32.5 3.5 94.1 45.5 70.1 76.2
+ DPO (Li et al., 2024) 4.1 56.7 54.3 2.2 95.0 58.8 73.1 81.5
+ TARS (Ours) 2.1 59.8 12.5 0.6 98.9 67.0 82.0 86.6
Appendix E Discussions and Insights
E.1 Why TARS outperforms DPO: Beyond numbers
While TARS consistently outperforms standard DPO across hallucination benchmarks, its effectiveness stems not only from empirical gains, but from the design principles that enable better preference alignment under uncertainty. Below, we outline the key factors contributing to the superior performance of TARS.
Token-level perturbation enhances alignment robustness. DPO relies on static textual inputs, making it susceptible to overfitting on superficial linguistic patterns such as high-frequency phrases or stylistic biases present in preference data. TARS addresses this issue by introducing controlled perturbations on visually agnostic tokens. These perturbations simulate semantically equivalent variations, thereby exposing the model to distributional shifts during training. As a result, the learned policy becomes more robust to alignment uncertainty and is encouraged to rely on visual grounding cues rather than memorized textual artifacts.
Visual-agnostic targeting preserves grounding fidelity. Unlike random or uniform perturbation strategies, TARS selectively perturbs tokens with low cross-modal relevance, those that carry minimal visual grounding. This design ensures that semantic shifts are injected without disrupting the causal connection between image and text. By isolating visually agnostic components for perturbation, TARS avoids damaging critical multimodal alignments, resulting in faithful responses that remain sensitive to visual semantics while being resilient to linguistic noise.
Spectral alignment encourages semantic consistency. To bridge the mismatch between input perturbations and static supervision, TARS introduces a spectral regularizer that aligns representations in the frequency domain. This global constraint allows for flexible modifications while maintaining semantic coherence at the sequence level. Unlike rigid token-level matching, frequency alignment smooths over local variations and discourages the model from latching onto spurious token-level correlations. This helps prevent overfitting to distribution-specific artifacts and improves generalization under preference supervision.
Table 7: Comparison of spectrum-based alignment (TARS) versus token-level contrastive alignment. Both models use identical perturbation policies. Spectrum-based alignment achieves lower hallucination and better semantic consistency.
Alignment Strategy AMBER MMHal POPE OBJHal
@@ E.1 Why TARS outperforms DPO: Beyond numbers
Token-level (contrastive) 3.4 57.2 16.3 1.4 2.36 0.49 87.1 93.3 15.8 4.9
Spectrum-based (TARS) 2.4 59.6 13.2 0.4 2.48 0.45 88.7 97.5 12.0 3.2
E.2 Token perturbation sensitivity
Impact of perturbation strength. TARS determines the number of perturbed tokens based on the model’s confidence and a scaling factor $\omega$. In Section C.1, we conduct a systematic evaluation of perturbation magnitude to assess sensitivity. Results show that moderate perturbation levels (e.g., $\omega=1\mathrm{e}{-3}$) yield optimal hallucination suppression, while excessively small or large values lead to under- or over-regularization. These findings confirm that the model benefits from controlled perturbation, and that TARS is robust across a range of $\omega$ values when properly calibrated.
Stability across perturbation strategies. We also compare two token-level perturbation strategies used in TARS: token masking and synonym replacement. Despite their distinct mechanisms—masking introduces structural noise, while replacement maintains fluent semantics—both consistently outperform unperturbed DPO. Notably, masking yields slightly better hallucination mitigation, likely due to its stronger distributional shift. The close performance of both strategies suggests that the core effectiveness of TARS is not overly sensitive to the specific perturbation operator, as long as semantic integrity is preserved.
Effectiveness of adaptive perturbation scope. Instead of applying a fixed number of perturbations, TARS dynamically scales the perturbation budget according to the model’s alignment uncertainty. This adaptive strategy ensures that confident predictions remain mostly intact, while uncertain ones are regularized more aggressively. Such input-aware perturbation improves training stability and avoids unnecessary semantic drift, reinforcing the model’s ability to distinguish visually grounded content from spurious correlations.
E.3 Spectrum-based alignment vs token-level alignment
Motivation. While TARS perturbs visually agnostic tokens to expose latent alignment vulnerabilities, the preference supervision signal remains static. This creates a mismatch: dynamic inputs versus fixed feedback. A common solution is to enforce token-level representation consistency, such as applying cosine similarity or contrastive loss between the perturbed and original hidden states. However, such approaches assume strict token-wise correspondence, which may not hold under semantic-preserving perturbations.
Limitations of token-level alignment. We empirically observe that enforcing rigid token-level matching often leads to instability during training and degrades hallucination mitigation. Token-level losses tend to penalize even small, semantically irrelevant shifts introduced by benign perturbations. This over-constraint reintroduces the spurious correlation that TARS aims to alleviate. In contrast, frequency-based alignment allows for local flexibility while enforcing global consistency in the hidden representation spectrum.
Comparative evaluation. To validate this claim, we compare the spectral preference loss in Equation (11) with a baseline token-level contrastive alignment strategy:
  • Contrastive loss (Token-level): Hidden states from $(x,q,y_{w})$ and $(x,\phi(q),y_{w})$ are aligned using an InfoNCE objective.
  • Spectrum loss (TARS): Frequency magnitude alignment via log-ratio of FFT-transformed hidden states.
As shown in Table 7, spectrum-based alignment achieves lower hallucination rates and better factual grounding across AMBER and OBJHal. This confirms that semantic-level flexibility, rather than rigid token consistency, is key to robust preference optimization.
E.4 Limitations
@@
In this work, we adopt two simple perturbation strategies: token masking and synonym replacement. These methods are chosen for their clarity, efficiency, and ease of interpretation, allowing us to isolate the effects of token-level alignment without introducing unnecessary complexity. However, their simplicity may limit the generality and flexibility of the approach. Future work could explore adaptive or data-driven perturbation mechanisms that better balance semantic preservation with distributional shift. Additionally, the current token selection strategy, based on cross-modal similarity heuristics, could be enhanced by learning-based relevance estimation or causal attribution techniques to enable more precise and effective perturbations.
Appendix F Qualitative Examples
@@
We provide qualitative comparisons between standard DPO and our proposed TARS in Table 8, across diverse image-question pairs. TARS consistently demonstrates improved grounding and hallucination suppression, outperforming traditional DPO in several key aspects:
@@
No degradation in response completeness. Importantly, TARS maintains response richness without sacrificing length or informativeness. As illustrated in (d), TARS generates a detailed yet grounded answer, whereas DPO provides longer responses but includes hallucinated attributes.
Better fine-grained grounding. TARS exhibits improved alignment at the fine-grained level, such as object color and count. In examples (b) and (e), it correctly identifies color attributes that DPO misrepresents.
Enhanced descriptive richness. TARS responses incorporate more relevant visual details, indicating stronger cross-modal alignment. The generated answers are not only more accurate but also semantically richer, reflecting a deeper understanding of the visual input.
These examples highlight the effectiveness of token-adaptive perturbation in guiding the model to learn robust visual-textual associations, ultimately reducing hallucinations without compromising informativeness.
Table 8: Qualitative comparisons between DPO and our proposed TARS across five diverse image-prompt pairs, denoted as (a)–(e). Each row presents the same visual input and accompanying question, with model responses shown for both methods. Hallucinated content is highlighted in red, while accurate visual grounding is marked in green. TARS consistently produces more faithful and informative responses, demonstrating superior grounding and hallucination mitigation even under visually ambiguous or linguistically subtle conditions.

### v4

@@ 10.1 AMBER Generative and Discriminative Metrics
We present extended results on the AMBER benchmark in Tab.˜9, evaluating hallucination performance from both generative and discriminative perspectives.
@@
Beyond generative evaluation, we further introduce fine-grained discriminative metrics that assess hallucination across four categories: Existence, Relation, Attribute, and Action. As shown in the right half of the table, TARS consistently outperforms both DPO and the LLaVA baseline in all dimensions.
Table 9: Comparison of generative and fine-grained discriminative hallucination metrics on the AMBER benchmark.
@@
LLaVA-7B [llava_origin] 7.9 54.7 37.1 3.2 82.9 58.6 65.6 70.1
+ DPO 4.9 56.6 26.4 2.5 87.1 59.7 74.6 79.4
+ TARS 2.4 59.6 13.2 0.4 95.3 62.8 78.6 86.5
LLaVA-13B [llava_origin] 6.7 52.1 32.5 3.5 94.1 45.5 70.1 76.2
+ DPO 4.1 56.7 54.3 2.2 95.0 58.8 73.1 81.5
+ TARS 2.1 59.8 12.5 0.6 98.9 67.0 82.0 86.6
10.2 Extended Hallucination Benchmarks
We provide extended evaluation on two additional benchmark suites, MS-COCO [lin2014microsoft] CHAIR and multimodal understanding benchmarks, to demonstrate that TARS improves hallucination suppression without degrading broader multimodal reasoning.
@@ 10.2 Extended Hallucination Benchmarks
LLaVA-Bench and SeedBench (Tab.˜11). To verify that hallucination suppression does not impair general multimodal capabilities, we evaluate on LLaVA-Bench [llava_origin] (conversation, detail, and reasoning sub-tasks) and SeedBench [li2023seedbench] (multi-choice visual reasoning). TARS improves over both the base LLaVA and DPO across all sub-tasks: the overall LLaVA-Bench score increases from 62.9 (LLaVA) to 67.2 (+4.3), and SeedBench accuracy rises from 32.4% to 38.7% (+6.3 pp). Notably, the largest gain appears in visual reasoning (+3.9 over LLaVA, +2.3 over DPO), indicating that TARS strengthens the model’s ability to perform multi-step inference grounded in visual evidence.
Table 10: Extended MS-COCO CHAIR evaluation across multiple settings.
MS-COCO CHAIR $_{I}^{-} CHAIR$$_{S}^{-}$ $^{1}_{I} CHAIR$$^{1}_{S}$ $^{2}_{I} CHAIR$$^{2}_{S}$
LLaVA-v1.5-7B 10.5 32.7 18.8 62.7 64.3 90.7
DPO 7.2 21.4 13.6 41.9 49.8 72.3
TARS (ours) 5.9 17.2 10.8 33.6 38.5 57.9
Table 11: Evaluation on LLaVA-Bench and SeedBench, showing TARS improves multimodal understanding.
Method Conv. Detail Reasoning Overall SeedBench Acc.
LLaVA-v1.5-7B 61.8 53.6 71.2 62.9 32.4
DPO 63.5 55.4 72.8 64.6 37.1
TARS (ours) 65.9 57.8 75.1 67.2 38.7
10.3 Comparison with Data Augmentation Baselines
A natural question is whether the gains of TARS can be replicated by simply augmenting the preference training data. We conduct a systematic comparison against two families of data augmentation strategies applied on top of standard DPO:
(1) Paraphrasing. We use GPT-3.5-Turbo to back-translate each preferred response through Chinese and French, then paraphrase back to English. This yields one additional variant per preference pair, preserving semantic content while diversifying surface forms. The augmented pairs are mixed with the original 4.8k samples during DPO training.
(2) LLM-based augmentation. We prompt an instruction-tuned LLM (Vicuna-13B [chiang2023vicuna]) to generate new preferred responses conditioned on the same image-question pairs, using the original preferred response as a reference. We experiment with two scales: $1\times$ (one new response per pair, yielding 9.6k total) and $5\times$ (five new responses per pair, yielding 28.8k total). Generated responses are filtered by CLIP score to remove low-quality samples before mixing with the original data.
Protocol. All models use LLaVA-v1.5-7B as the base and train with identical hyperparameters (learning rate, batch size, number of epochs) to isolate the effect of data composition. The only variable is the training data: original 4.8k samples vs. augmented sets.
Analysis. As shown in Tab.˜12, augmentation provides incremental improvements: paraphrasing reduces Hal-Rate from 26.4% to 24.8% (–1.6 pp), and $5\times$ LLM augmentation further lowers it to 16.0% (–10.4 pp). However, even the most aggressive augmentation ($5\times$, using $6\times$ more data) still underperforms TARS on every metric (Hal-Rate 16.0% vs. 13.2%, Cog 1.2 vs. 0.4). This gap demonstrates that TARS provides a fundamentally different and stronger form of robustness that cannot be recovered through data scaling alone.
Crucially, TARS and augmentation are complementary: combining TARS with $5\times$ LLM augmentation achieves the best overall performance (CHAIR 2.1, Cover 60.8%, Hal-Rate 12.3%). This suggests that augmentation diversifies the preference distribution while TARS sharpens the model’s sensitivity to token-level hallucination cues, and the two mechanisms address orthogonal failure modes.
Table 12: Comparison of TARS against data augmentation baselines across four hallucination benchmarks. Paraphrasing denotes back-translation-based response diversification; LLM Aug. denotes LLM-generated preferred responses, where $1\times$ and $5\times$ indicate the expansion factor relative to the original 4.8k preference set. All methods build on standard DPO. TARS is complementary to augmentation strategies.
Method AMBER MMHal POPE OBJHal
[…]
LLaVA-v1.5-7B-DPO 4.9 56.6 26.4 2.5 2.19 0.61 87.8 82.0 14.0 5.0
+ Paraphrasing 4.5 57.9 24.8 2.3 2.24 0.58 87.9 83.5 13.6 4.8
[…]
+ TARS 2.4 59.6 13.2 0.4 2.48 0.45 88.7 97.5 12.0 3.2
+ TARS & Paraphrasing 2.3 60.1 12.8 0.4 2.50 0.44 88.8 97.6 11.5 3.1
[…]
10.4 Comparison with Retrieval-Augmented DPO
RE-ALIGN [xing2025realign] augments standard DPO by injecting retrieved visual evidence, specifically relevant image patches and captions retrieved via CLIP, into the preference optimization pipeline. This approach enhances the factual grounding of preference pairs by providing the model with additional visual context during training. In contrast, TARS improves robustness through an entirely orthogonal mechanism: adversarial token-level perturbation within a min–max framework, without modifying the input data or retrieval pipeline.
Experimental setup. We compare four configurations on LLaVA-v1.5-7B using the same 4.8k preference samples: base LLaVA, standard DPO, DPO + RE-ALIGN, and TARS + RE-ALIGN. For RE-ALIGN, we follow the original implementation using CLIP-ViT-L/14 for retrieval and inject the top-3 retrieved evidence snippets.
Analysis. As shown in Tab.˜13, RE-ALIGN alone provides strong improvements over standard DPO, reducing Hal-Rate from 26.4% to 16.8% (–9.6 pp) and improving object coverage from 56.6% to 58.9% (+2.3 pp). However, combining TARS with RE-ALIGN yields substantial further gains across all six metrics: Hal-Rate drops to 12.5% (–4.3 pp below RE-ALIGN alone), CHAIR decreases from 3.6 to 2.2, and MMH improves from 2.31 to 2.57.
The complementarity of the two methods can be understood through their distinct mechanisms: RE-ALIGN enriches the input space by providing additional visual evidence, reducing hallucinations caused by insufficient grounding information. TARS, on the other hand, strengthens the optimization process by exposing the model to adversarial token perturbations during training, reducing hallucinations caused by fragile token-level associations. Together, they address both data-side and algorithm-side sources of hallucination, achieving the best performance across all metrics.
Table 13: Complementarity of TARS and RE-ALIGN. Combining both methods achieves the best performance.
[…]
LLaVA-v1.5-7B 7.6 51.7 35.4 4.2 2.02 0.61
+ DPO 4.9 56.6 26.4 2.5 2.19 0.61
+ RE-ALIGN 3.6 58.9 16.8 1.3 2.31 0.52
+ TARS & RE-ALIGN 2.2 60.4 12.5 0.4 2.57 0.43
11 Additional Results on Muffin
We further validate TARS on the Muffin-13B architecture (Tab.˜14). Consistent with our findings on LLaVA, both perturbation strategies yield substantial improvements over DPO and CHiP-DPO. TARS with token masking achieves the strongest overall performance, while synonym replacement remains competitive. These results confirm the versatility of our approach across different MLLM backbones.
Table 14: Comparison of hallucination benchmarks on Muffin-13B. Bold denotes the best, and underlined denotes the second-best.
Algorithm AMBER MMHal POPE OBJHal
[…]
Muffin-13B [yu2023reformulating] 7.5 45.7 34.6 3.4 2.27 0.58 83.0 80.7 47.3 15.2
+ RLHF [RLHF] 7.1 45.2 37.1 3.5 2.12 0.64 84.0 79.8 45.5 12.7
+ DPO [li2024multi] 6.0 46.4 29.6 2.8 2.45 0.55 83.7 81.2 43.8 13.9
+ CHiP-DPO [CHIP] 4.8 48.2 18.9 1.7 2.70 0.47 84.5 82.1 35.2 11.5
+ TARS (Mask) 3.6 49.5 16.2 1.4 2.75 0.41 87.4 84.9 28.7 8.2
+ TARS (Replace) 4.0 48.9 16.5 1.5 2.76 0.43 86.5 83.8 29.3 8.8
12 Discussions and Insights
12.1 Why TARS Outperforms DPO: Beyond Numbers
While TARS consistently outperforms standard DPO across hallucination benchmarks, its effectiveness stems not only from empirical gains, but from the design principles that enable better preference alignment under uncertainty.
Token-level perturbation enhances alignment robustness. DPO relies on static textual inputs, making it susceptible to overfitting on superficial linguistic patterns. TARS addresses this issue by introducing controlled perturbations on visually agnostic tokens. These perturbations simulate semantically equivalent variations, exposing the model to distributional shifts during training. As a result, the learned policy becomes more robust to alignment uncertainty.
Visual-agnostic targeting preserves grounding fidelity. Unlike random or uniform perturbation strategies, TARS selectively perturbs tokens with low cross-modal relevance. This design ensures that semantic shifts are injected without disrupting the causal connection between image and text, resulting in faithful responses that remain sensitive to visual semantics while being resilient to linguistic noise.
Spectral alignment encourages semantic consistency. TARS introduces a spectral regularizer that aligns representations in the frequency domain. This global constraint allows for flexible modifications while maintaining semantic coherence at the sequence level, discouraging the model from latching onto spurious token-level correlations.
Table 15: Comparison of spectrum-based alignment (TARS) versus token-level contrastive alignment. Both models use identical perturbation policies.
Strategy AMBER MMHal POPE OBJHal
[…]
Token-level 3.4 57.2 16.3 1.4 2.36 0.49 87.1 93.3 15.8 4.9
Spectrum (TARS) 2.4 59.6 13.2 0.4 2.48 0.45 88.7 97.5 12.0 3.2
12.2 Limitations
@@
In this work, we adopt two simple perturbation strategies: token masking and synonym replacement. These methods are chosen for their clarity, efficiency, and ease of interpretation. However, their simplicity may limit the generality and flexibility of the approach. Future work could explore adaptive or data-driven perturbation mechanisms [bai2026dice] that better balance semantic preservation with distributional shift. Additionally, the current token selection strategy, based on cross-modal similarity heuristics, could be enhanced by learning-based relevance estimation or causal attribution techniques.
13 Qualitative Examples
@@
We provide qualitative comparisons between standard DPO and our proposed TARS in Fig.˜8, across diverse image-question pairs. TARS consistently demonstrates improved grounding and hallucination suppression, outperforming traditional DPO in several key aspects:
@@
No degradation in response completeness. TARS maintains response richness without sacrificing length or informativeness.
Better fine-grained grounding. TARS exhibits improved alignment at the fine-grained level, such as object color and count.
Enhanced descriptive richness. TARS responses incorporate more relevant visual details, indicating stronger cross-modal alignment.
Figure 8: Qualitative comparisons between DPO and TARS across five diverse image-prompt pairs, denoted as (a)–(e). Hallucinated content is highlighted in red, while accurate visual grounding is marked in green. TARS consistently produces more faithful and informative responses.